Maxwell Manufacturing issued $750,000, 10-year, 10% bonds at 105.

What is the total amount of interest expense that will be recorded over the life of these bonds?

Multiple Choice

  • $825,000

  • $787,500

  • $750,000

  • $712,500
Answer #1

Premium on bonds = 750000*1.05 - 750000 = $37500

Interest expense over the life of the bonds = 750000*10%*10 - 37500 = $712500